If the rms velocity of a gas is 100m/sec. Calculate its average velocity.

Grade:12th pass

1 Answers

Arun
6 years ago
We know that
Vrms = \sqrt(3RT/M)
Vavg = \sqrt(8RT/\pie\LARGE \piM)
Hence
Vavg/ Vrms = \LARGE \sqrt(8/ 3\LARGE \pi)
So
Vavg = Vrms * 0.9213
Vavg = 100 * 0.9213= 92.13 m/sec